OpenStudy (anonymous):

Ashley has scored 30, 23, and 25points in her three basketball games so far. How many points does she need to score in her next game so that her average (mean) is 24 points per game?

OpenStudy (anonymous):

easy!

OpenStudy (anonymous):

ok now let me guide you through this

OpenStudy (anonymous):

I need the explain

OpenStudy (anonymous):

it says she should get an average of 24 points per game right .. that means she got 24 points in each game ... let me show by a diagram

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[(30+23+25+x)/4=24 points\]

OpenStudy (anonymous):

do you agree with the formula i wrote?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok good .. now see just transfer the four to the other side

OpenStudy (anonymous):

you get ...

OpenStudy (anonymous):

\[30+25+23+x=24*4=>96\]

OpenStudy (anonymous):

just equate now ... bring the 30 25 and 24 to the other side you get 78 therefore x=18

OpenStudy (anonymous):

Now I understand .. Thank you ,, you are good explainer :)

OpenStudy (anonymous):

dont forget to fan ! cheers glad to help you!

OpenStudy (anonymous):

Ok =)